a student takes an education loan from a bank. He expects to borrow $40,000 a year for the next 4 years (the first amount is drawn immediately, on 2/28/2010). The annual interest rate being charged is 8%; for the first several years, the interest is not paid, but it is added to the loan at the end of each year. He is required to repay the loan in equal annual payments starting 7 years from the date of the loan (i.e., the first payment will be made on 2/28/2017), with one, payment per year for 15 years.

a. What will be the amount of each loan payment?

b. Right after making the fifth annual payment, he decides to repay his entire loan. How much will he have to pay?

c. How much interest will he end up paying?
Answer #1

1.
=40000/8%*(1.08^4-1)*1.08=194664.0384

2.
=194664.0384*1.08^4*(1-1/1.08)/(1-1/1.08^15)=28649.01404

3.
=(194664.0384*1.08^3*1.08^5)-28649.01404/0.08*(1.08^5-1)=192237.2163

4.
=28649.01404*15-40000*4=269735.2106